OpenStudy (anonymous):

Juan and Nina are in the Junior High Orchestra. As part of the orchestra they each have to learn two instruments. Juan is learning to play the violin and the cello. For every 3 hours that Juan spends practicing the violin he practices the cello for 2.5 hours. Nina is learning to play the viola and cello. For every 3 hours Nina spends practicing the viola she practices the cello for 2 hours. If Juan practiced the violin for 9 hours last week and Juan and Nina practiced the cello for the same amount of time, then Nina practiced the viola for____ hours last week.

OpenStudy (anonymous):

Answer choices please?

OpenStudy (triciaal):

Juan 3 V = 2.5 C when V = 9 then C = 3 * 9 /2.5 then same amount of time Nina practiced the C Nina 3 V = 2 C V = 2 C /3 = 2 (C above) / 3 V = violin for Juan and Viola for Nina, C = Cello

OpenStudy (anonymous):

J=3/2.5 N=3/2 J practices violin for 9 hours J practices cello for 22.5